# Gamma-ray burst pulse structures and emission mechanisms

### v1

@@ 3.3 Relation between the spectral shape and pulse shape
Correlation analysis shows that Spearman’s, when applied to the full sample, is $r_{s}=0.23$, with p-value = $0.07$. This indicates a trend between the maximal value of the low energy spectral index, $\alpha_{\max}$ and the pulse shapes: the more FRED-like pulses tend to have softer $\alpha_{\max}$ values. This may suggest a non-thermal emission origin. Generally, one finds that the more symmetric-like pulses have steeper spectral slopes (larger values of $\alpha_{\max}$), which can be interpreted as having a thermal (or quasi-thermal) origin.
The lower $\alpha_{\max}$ values of FRED-like pulses may point to the idea that they originate at a large radius in the jet; hence, their spectra appear non-thermal, and the asymmetry in the light curve is due to light aberration. On the other hand, the higher value of $\alpha_{\max}$ of symmetric-like pulses hints towards the possibility that they originate at smaller radii, where the photospheric component is more pronounced, and photon diffusion below the photosphere symmetrize the pulse. There is also no substantial difference between the single peak and combined peak pulse results, as we can see from Spearman’s $r_{s}$ values of $0.24$ and $0.20$ respectively.
@@ 4.1 Pulse symmetry
The pulse analysis method introduced in subsection 2.2 has been successful in quantifying the temporal symmetry observed in light curves. We are unaware of any instrumental bias capable of causing small amplitude deviations in a FRED curve that would mimic the bright, long features seen as a symmetric-like curve. From Figure 4, one sees that many of the pulses in our sample are symmetric-like and that their fraction is particularly high for the initial pulses in a given burst. We acknowledge that Hakkila et al. (2018) previously identified temporal symmetry in some BATSE light curves, although this work did not exclusively consider multi-pulses.
The finding that a large fraction of the observed pulses in the prompt GRB light curve are symmetric is surprising. The prompt emission is produced as a result of energy dissipation processes that occur in relativistic jets. The rise time scales of the pulse indicate a continuous dissipation of energy: as more and more energy is dissipated, more and more photons are emitted. On the other hand, the decay is subject to various light aberration effects, such as high-latitude emission. There is, therefore, no apriori reason to obtain a symmetric pulse light curve, as is observed. In order to address this problem, Hakkila & Nemiroff (2019) suggested that some aspects of the pulsed emission may be repeated in reverse order as a result of events in the jet; either the emission mechanism must take this into account through the jet’s kinematic behaviour or a different emission mechanism is required to explain the time-reversed portion of the light curve.
@@
The majority of the pulses in our sample lie between the slow-cooling synchrotron and the NDP line, above which only a photospheric component is theoretically expected. Intermediate values can, therefore, be explained as a simultaneous contribution from both thermal (photospheric) and non-thermal (synchrotron) emission. Harder $\alpha_{\max}$ values of the initial symmetric-like pulses in a given burst can be explained as a stronger contribution of thermal emission components in the overall spectra. The more FRED-like pulses have soft $\alpha_{\max}$ values indicating non-thermal emission origin. This result is a natural outcome of the classical GRB “fireball” model: thermal emission originates from the photosphere, which is the closest radius to the progenitor from which photons escape and is thus expected to be the first signal detected. Non-thermal spectra are possibly emitted from larger radii (well above the photosphere), resulting in more FRED-like light curves. We conclude that a different emission radius is the most likely origin of the various pulse shapes.
@@ 4.2 Pulse duration effects on the shape
From the results presented in Figure 9, we conclude that the pulse duration does not correlate with the pulse shape and, by extension, the pulse emission mechanism. This seems counter-intuitive since the pulse durations in our sample vary by one order of magnitude (from a few seconds to $>50$ s). This indicates a similarity between different pulses, hinting towards a common underlying origin. One caveat is that the redshifts of most GRBs in our sample (all except four) are unknown. Therefore, we could not correct for the different redshifts.
@@ 4.3 On the novelty of the fitting function
As a fitting function, we consider two approximate sigmoid functions which independently fit for the rising slope $s_{l}$ and decaying slope $s_{r}$. The half-time of the rise phase (pulse start to peak) and the decay phase (peak to pulse end) is represented by $r_{l}$ and $r_{r}$, respectively (see Equation 1). In our function, these are treated as independent parameters. This is the key difference from previous fits (Norris et al., 1996; Hintze, 2022), which enables the provision of good fits to wide pulses whose peak extends over a substantial duration. This additional degree of freedom is crucial in obtaining symmetric pulse shapes. For optimal fits, we always require $r_{l}\leq r_{r}$; otherwise, the obtained slope values do not provide insight into the observed ones. This condition imposes a restriction on the fits, which, however, do not affect the overall conclusion.
@@ 4.4 Theoretical interpretation
Here, $L$ is the GRB luminosity, $\Gamma$ is the jet Lorentz factor, $\sigma_{T}$ is the Thomson cross section, $m_{p}$ is the proton mass, and $c$ is the speed of light. For typical values of outflow parameters, $\Gamma\gtrsim 10^{2}$, the photospheric radius is in the range $10^{11}-10^{13}$  cm, and the resulting timescale is only a fraction of a second. This timescale is shorter by 2-4 orders of magnitude than the typical pulse duration in our sample, which is about a second or more. However, this time scale is extremely sensitive to the uncertain value of the jet Lorentz factor, as it is proportional to $\Gamma^{-5}$. Therefore, if the typical Lorentz factor of GRBs in our sample is lower than the commonly assumed value and is rather of the order of only a few tens, as recently suggested (Dereli-Bégué et al., 2022), this discrepancy disappears. The results presented here may, therefore, provide additional, indirect evidence that the Lorentz factor of many GRBs may be only a few tens rather than a few hundreds, as is commonly assumed. For further indications on low values of GRB Lorentz factors, see Dereli-Bégué et. al., ( 2024, in prep.)
@@ Acknowledgments
GA thanks Dhruva Sambrani and Damien Bégué for the helpful discussion and comments. This research made use of the High Energy Astrophysics Science Archive Research Center Online Service HEASARC at the NASA/Goddard Space Flight Center. AP acknowledges the support from the European Union (EU) via ERC consolidator grant 773062 (O.M.J.), and from the Israel Space Agency via grant #6766. FR acknowledges support from the Swedish National Space Agency (2021-00180 and 2022-00205).

### v2

@@ 3.3 Relation between the spectral shape and pulse shape
Correlation analysis shows that Spearman’s, when applied to the full sample, is $r_{s}=0.23$, with p-value = $0.07$. This indicates a trend between the maximal value of the low energy spectral index, $\alpha_{\max}$, and the pulse shapes: the more FRED-like pulses tend to have softer $\alpha_{\max}$ values. This may suggest a non-thermal emission origin. Generally, one finds that the more symmetric-like pulses have steeper spectral slopes (larger values of $\alpha_{\max}$), which can be interpreted as having a thermal (or quasi-thermal) origin.
The lower $\alpha_{\max}$ values of FRED-like pulses may point to the idea that they originate at a large radius in the jet; hence, their spectra appear non-thermal, and the asymmetry in the light curve is due to light aberration. On the other hand, the higher value of $\alpha_{\max}$ of symmetric-like pulses hints towards the possibility that they originate at smaller radii, where the photospheric component is more pronounced, and photon diffusion below the photosphere symmetrizes the pulse. There is also no substantial difference between the single peak and combined peak pulse results, as we can see from Spearman’s $r_{s}$ values of $0.24$ and $0.20$ respectively.
@@ 4.1 Pulse symmetry
The pulse analysis method introduced in subsection 2.2 has been successful in quantifying the temporal symmetry observed in light curves. We are unaware of any instrumental bias capable of causing small amplitude deviations in a FRED curve that would mimic the bright, long features seen as a symmetric-like curve. From Figure 4, one sees that many of the pulses in our sample are symmetric-like and that their fraction is particularly high for the initial pulses in a given burst.
The finding that a large fraction of the observed pulses in the prompt GRB light curve are symmetric is surprising. The prompt emission is produced as a result of energy dissipation processes that occur in relativistic jets. The rise time scales of the pulse indicate a continuous dissipation of energy: as more and more energy is dissipated, more and more photons are emitted. On the other hand, the decay is subject to various light aberration effects, such as high-latitude emission. There is, therefore, no apriori reason to obtain a symmetric pulse light curve, as is observed.
Several authors (Hakkila et al., 2018; Hakkila & Nemiroff, 2019; Hakkila, 2021; Liu & Zou, 2023) applied a fitting approach that combines a monotonic component with fast-varying residual structures. They found that these residuals exhibit temporal symmetry, with patterns in the decay phase appearing as time-reversed counterparts of those in the rise phase, typically reflected about the pulse peak time. In our sample of GRB lightcurves, we also identify time-reversed structures in the rising and decay phases of the pulses. We account for these underlying variations in the pulse definition itself and by the inclusion of the combined peak pulse category.
To explain the symmetric nature of these structures, Hakkila & Nemiroff (2019) suggested that some aspects of the pulsed emission may be repeated in reverse order as a result of events in the jet; either the emission mechanism must take this into account through the jet’s kinematic behaviour or a different emission mechanism is required to explain the time-reversed portion of the light curve.
@@
The majority of the pulses in our sample lie between the slow-cooling synchrotron and the NDP line, above which only a photospheric component is theoretically expected. Intermediate values can, therefore, be explained as a simultaneous contribution from both thermal (photospheric) and non-thermal (synchrotron) emission. Harder $\alpha_{\max}$ values of the initial symmetric-like pulses in a given burst can be explained as a stronger contribution of thermal emission components in the overall spectra. The more FRED-like pulses have soft $\alpha_{\max}$ values, indicating non-thermal emission origin. This result is a natural outcome of the classical GRB “fireball” model: thermal emission originates from the photosphere, which is the closest radius to the progenitor from which photons escape and is thus expected to be the first signal detected. Non-thermal spectra are possibly emitted from larger radii (well above the photosphere), resulting in more FRED-like light curves. We conclude that a different emission radius is the most likely origin of the various pulse shapes.
@@ 4.2 Pulse duration effects on the shape
From the results presented in Figure 9, we conclude that the pulse duration does not correlate with the pulse shape and, by extension, the pulse emission mechanism. This seems counterintuitive since the pulse durations in our sample vary by one order of magnitude (from a few seconds to $>50$ s). This indicates a similarity between different pulses, hinting towards a common underlying origin. One caveat is that the redshifts of most GRBs in our sample (all except four) are unknown. Therefore, we could not correct for the different redshifts.
@@
Initially, results from prompt emission pulses observed by BATSE showed that the pulse width remains constant throughout the GRB time history (Ramirez-Ruiz & Fenimore, 2000). However, using our Fermi-GBM dataset, we now show that later pulses tend to be longer in duration than earlier ones, with a Spearman correlation coefficient $r_{s}$ = 0.25 observed between the pulse duration and pulse number. These findings are consistent with the results reported by Hakkila et al. (2024). This, along with the hard-to-soft evolution, appears consistent with the known properties of X-ray flares (Chincarini et al., 2007, 2010; Margutti et al., 2011; Dereli-Bégué et al., 2025).
@@ 4.3 On the novelty of the fitting function
As a fitting function, we consider two approximate sigmoid functions that independently fit for the rising slope $s_{l}$ and decaying slope $s_{r}$. The half-time of the rise phase (pulse start to peak) and the decay phase (peak to pulse end) is represented by $r_{l}$ and $r_{r}$, respectively (see Equation 1). In our function, these are treated as independent parameters. This is the key difference from previous fits (Norris et al., 1996; Hintze, 2022; Norris et al., 2005), which enables the provision of good fits to wide pulses whose peak extends over a substantial duration. This additional degree of freedom is crucial in obtaining symmetric pulse shapes. For optimal fits, we always require $r_{l}\leq r_{r}$; otherwise, the obtained slope values do not provide insight into the observed ones. This condition imposes a restriction on the fits, which, however, do not affect the overall conclusion.
Due to our pulse definition criteria, we managed to extract a large number of pulses from these bursts compared to previous studies. The pulse definition we use is purely dependent on the S/N ratio of the count rates. This definition may differ from the criteria that Hakkila & Preece (2014); Hakkila et al. (2018) and Hakkila (2021) applied to the BATSE pulse identification, which further considers the residual structures along with the monotonic pulse fit. It is also possible that the $2\mu s$ resolution for the GBM TTE data might be influencing the pulse identification because of the increased sensitivity to variations, causing larger pulse numbers. These factors might be the cause of the difference in the number of pulses found in the two analyses.
While previous studies have made the delineation between pulse and structures (Hakkila, 2021; Liu & Zou, 2023), there is not yet a consensus on a physical model of the emission mechanism requiring these structures. Hence, here we pursue an alternative, empirical perspective of pulses, which is mathematically well-defined and reproducible without imposing any assumption of pulse morphologies on the data. Due to the flexible approach of the pulse definition used in this work, we are able to capture a broader range of pulse shapes for both single and multipulse GRBs, which can include variations in the pulse shapes. This offers a complementary approach to previous light curve studies.
We acknowledge that the flexibilities in our pulse definition allow for a broader range of pulse morphologies. In future studies, it will be interesting to compare how the correlation results obtained from different pulse definitions depend on the initial selection criteria. Although we emphasize that the trends we see in this study are self-consistent within the framework we use.
@@ 4.4 Theoretical interpretation
Here, $L$ is the GRB luminosity, $\Gamma$ is the jet Lorentz factor, $\sigma_{T}$ is the Thomson cross section, $m_{p}$ is the proton mass, and $c$ is the speed of light. For typical values of outflow parameters, $\Gamma\gtrsim 10^{2}$, the photospheric radius is in the range $10^{11}-10^{13}$  cm, and the resulting timescale is only a fraction of a second. This timescale is shorter by 2-4 orders of magnitude than the typical pulse duration in our sample, which is about a second or more. However, this time scale is extremely sensitive to the uncertain value of the jet Lorentz factor, as it is proportional to $\Gamma^{-5}$. Therefore, if the typical Lorentz factor of GRBs in our sample is lower than the commonly assumed value and is rather of the order of only a few tens, as recently suggested (Dereli-Bégué et al., 2022), this discrepancy disappears. The results presented here may, therefore, provide additional, indirect evidence that the Lorentz factor of many GRBs may be only a few tens rather than a few hundreds, as is commonly assumed. For further indications on low values of GRB Lorentz factors, see Dereli-Bégué et al. (2025).
@@ Acknowledgments
GA thanks Dhruva Sambrani and Damien Bégué for the helpful discussion and comments. We also thank the anonymous referee for suggestions that improved the clarity of the manuscript. This research made use of the High Energy Astrophysics Science Archive Research Center Online Service HEASARC at the NASA/Goddard Space Flight Center. AP acknowledges the support from the European Union (EU) via ERC consolidator grant 773062 (O.M.J.), and from the Israel Space Agency via grant #6766. FR acknowledges support from the Swedish National Space Agency (2021-00180 and 2022-00205).
Appendix A Comparison of the fitting function used here and in Norris et al. (2005)
Here we compare the fits carried out using Equation 1 with fitting functions that appear previously in the literature, specifically the one used in Norris et al. (2005). We applied the Norris et al. (2005) model to our sample of 61 pulses from 22 GRBs. To ensure consistency, we have applied the same sample selection processes to all GRB pulses. To demonstrate the advantage of Equation (1), we give here two examples of pulse shape fits.
Figure A.1: Comparing fit of GRB190114 using the Norris et. al. (2005) function [left] and the function used here [right]. The solid green line represents the raw count-rate data from GBM, and the dashed blue line represents the pulse shape model fit according to Norris et al. (2005) Function and Equation 1, respectively.
Figure A.2: Comparing fit of GRB 120129 using the Norris et. al. (2005) function [left] and our function used here [right]. The solid green line represents the raw count-rate data from GBM, and the dashed blue line represents the pulse shape model fit according to Norris et al. (2005) Function and Equation 1, respectively.
In Figure A.1 we present the example of GRB 190114 fitted twice: once with the function used by Norris et al. (2005) [left] and using Equation (1) [right]. This lightcurve shows a steep decay, which is better captured using our Equation (1): The Norris et al. (2005) function fit gives $k=0.53\pm 0.02$ while we obtain $\phi=2.12\pm 0.22$. The $r^{2}$ values are similar - 0.89 and 0.94 for the two fits, respectively. As is demonstrated in Figure A.1, Equation (1) is more flexible in fitting pulses with longer rise times compared to decay times.
A second example is given by fitting the pulse of GRB 120129, shown in Figure A.2. While both the Norris et al. (2005) function [left] and Equation (1) [right] provide good fits to the pulse shape - both give $r^{2}=0.96$, we point out that the Norris function fit gives $k=0.57\pm 0.03$, which is the same value as obtained for fitting the pulse of GRB 190114 shown in Figure A.1 above; while in the fit carried using our Equation (1), we obtain $\phi=0.51\pm 0.06$, which is different than the value obtained above, as is evident from the pulse shape. Due to the wider range of possible $\phi$ values compared to the $k$ values, the fits of Equation (1) can provide more distinctive values for each pulse shape.
Furthermore, applying the Norris et al. (2005) function to our sample set, we found that 12 pulses did not meet the goodness-of-fit criteria $r^{2}>0.7$. We further found that 7 pulse shape fits were slightly better with the Norris et al. (2005) model compared with Equation (1). However, the maximum improvement in goodness-of-fit $r^{2}$ was only $\sim 2.75\%$. For example, the Norris et al. (2005) function fit of GRB 131014 provides a fit with $r^{2}=0.92$, whereas the function we use gives $r^{2}=0.90$. In all other pulses, Equation (1) produced better fits.